OpenStudy (anonymous):

use elimination to solve the system of equations given by 3x-2y=12 and 3x+5y=33

OpenStudy (anonymous):

Do you know which variable you can eliminate first?

OpenStudy (anonymous):

would it be x?

OpenStudy (anonymous):

Yup! 3x-3x=0

OpenStudy (anonymous):

3x-2y=12 3x+5y=33 ---------- -7y=-21 y=?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Yup! Now plug the 3 in any of the equations to find x 3x-2(3)=12 OR 3x+5(3)=33

OpenStudy (anonymous):

so would x equal 6?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Just kidding yea that's right

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Anytime